# Two Heads Are Better Than One: A Multi-Agent System Has the Potential to Improve Scientific Idea Generation

### v1

@@ Abstract
The rapid advancement of scientific progress requires innovative tools that can accelerate discovery. While recent AI methods, particularly large language models (LLMs), have shown promise in tasks such as hypothesis generation and experimental design, they fall short in replicating the collaborative nature of real-world scientific practices, where diverse teams of experts work together to tackle complex problems. To address the limitation, we propose an LLM-based multi-agent system, i.e., Virtual Scientists (VirSci), designed to mimic the teamwork inherent in scientific research. VirSci organizes a team of agents to collaboratively generate, evaluate, and refine research ideas. Through comprehensive experiments, we demonstrate that this multi-agent approach outperforms the state-of-the-art method in producing novel and impactful scientific ideas, showing potential in aligning with key insights in the Science of Science field. Our findings suggest that integrating collaborative agents can lead to more innovative scientific outputs, offering a robust system for autonomous scientific discovery.
@@ 1 Introduction
The rapid acceleration of scientific progress necessitates more efficient and innovative tools for exploring new concepts and tackling complex challenges (Park et al., 2023b). The concept of automatic scientific discovery has emerged as a promising solution to expedite innovation, representing a long-standing ultimate goal within the scientific community (Langley, 1987). With the development of artificial intelligence (AI), automatic scientific discovery has witnessed the potential to revolutionize how research is conducted by automating key steps in the scientific process, ranging from hypothesis generation to experimental design (Raghu & Schmidt, 2020; Spangler et al., 2014).
@@
Figure 1: The proposed LLM-based multi-agent system, VirSci, consists of five key steps: Collaborator Selection, where a research team is assembled; Topic Discussion, where the research topic is determined; Idea Generation, where team members propose and refine ideas; Novelty Assessment, where ideas are evaluated and voted on to select the best one; and Abstract Generation, where the selected idea is developed into a complete abstract.
More recently, foundational models, especially the large language models (LLMs) (OpenAI, 2023; Dubey et al., 2024), have shown significant progress in general capabilities, facilitating their applications in various stages of scientific discovery (Abramson et al., 2024; Chang & Ye, 2024), including literature reviews (Hsu et al., 2024), experimental designs (Huang et al., 2024), etc. A notable development is the AI Scientist (Lu et al., 2024), which introduces a scalable system for end-to-end scientific paper generation, highlighting the potential of LLMs to drive autonomous scientific discovery. Despite its capabilities, AI Scientist operates with ONE agent, which falls short of replicating real-world scientific practices, where research is often conducted through collaborative efforts involving diverse teams of experts (Kayacik et al., 2019).
To address the limitations of a single executive system, such as the AI Scientist, and to better replicate the collaborative nature of real-world scientific discovery (Gauch, 2003), we focus on the idea generation phase in the research process, which demonstrates more collaborative aspects (Linsey et al., 2005). From this end, we propose an LLM-based multi-agent system, Virtual Scientists (VirSci), designed to harness the potential of LLM agents in assisting autonomous scientific idea generation. Leveraging the inherent human-like reasoning capabilities of LLMs (Xie et al., 2024), VirSci simulates the collaborative process of research idea generation, which is divided into five steps (see Fig. 1): (1) Collaborator Selection, (2) Topic Selection, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. To be more specific, we construct a knowledge bank of the background of scientists of interest and develop digital twin (Tao et al., 2018) agents of real scientists using a retrieval-augmented generation (RAG) framework (Gao et al., 2023). The lead agent, or team leader, identifies appropriate collaborators based on the scientist collaboration network, aligning their expertise and research interests to reflect real-world cooperation patterns. Using a past paper database, the team retrieves reference works to guide the formation of novel ideas. Moreover, we implement a “team discussion” mechanism in subsequent stages, where collaborators engage in iterative inter- and intra-refinement dialogues, to enhance the quality of each output. Finally, the team generates a comprehensive abstract representing the proposed ideas. To evaluate the novelty of the ideas, we introduce a benchmark for the measurement from three perspectives: dissimilarity to past papers, alignment with contemporary research trends, and the potential influence in contemporary research (Shao et al., 2020; Yang et al., 2022). By comparing the abstract against both a past and a contemporary paper database, we ensure the generated ideas are innovative while aligning with emerging scientific directions, thus validating the effectiveness of our approach.
We conduct comprehensive experiments to verify the effectiveness of VirSci in producing novel scientific ideas. The findings demonstrate that the multi-agent system outperforms the single-agent executive with an average gain of 13.8% and 44.1% in alignment level and potential impact on contemporary research, respectively. Additionally, our experiments reveal emergent social behaviors among agents, aligning with prior studies in Science of Science (Fortunato et al., 2018; Wu et al., 2019; Zeng et al., 2021; Shi & Evans, 2023), suggesting the potential for further exploration of the mechanisms in research collaboration using multi-agent simulations. To sum up, our core contributions are summarized as follows:
@@
  • To the best of our knowledge, we propose the first multi-agent system for conducting scientific collaborations in an end-to-end pipeline from team organization to novel scientific idea generation. Furthermore, the real data is utilized for role-play and the objective evaluation of final outputs.
  • We conduct extensive evaluations to investigate VirSci in terms of the team settings and the novelty of generated scientific ideas. The results demonstrate that multi-agent collaboration can improve the quality of the outcomes, surpassing the SOTA single-agent method.
  • The simulation results align with the important findings in Science of Science, such as fresh teams tend to create more innovative research, showcasing the potential of VirSci as a powerful tool for future research in this field.
@@ 2.1 AI for Scientific Discovery
In recent years, AI has fundamentally reshaped the landscape of scientific discovery by providing powerful tools that enhance various research processes (Xu et al., 2021). AI techniques, especially generative AI, can facilitate basic scientific discoveries, such as identifying complex molecular (Vignac et al., 2022) and protein structures (Abramson et al., 2024), drastically reducing the time required for experimental iterations. These advancements have found wide application across diverse fields such as chemistry (Liu et al., 2023a), meteorology (Bi et al., 2023), and medicine (Rajpurkar et al., 2022), etc. Besides, with the advent of LLMs, AI methodologies can step further and collaborate in streamlining critical stages of the scientific pipeline, including hypothesis generation, experimental design, data acquisition, and analysis(Zheng et al., 2023; Wang et al., 2023; Miret & Krishnan, 2024; Wysocki et al., 2024; Lu et al., 2024). Nevertheless, these approaches lack the collaborative nature of the scientists intrinsic to real-world research. VirSci is the first to harness the power of an LLM-based multi-agent system to facilitate the generation of research ideas in autonomous scientific discovery.
2.2 Multi-agent Systems in Team Collaboration
A multi-agent system for team collaboration leverages autonomous agents to coordinate, communicate, and solve tasks within a shared environment, mimicking human teamwork dynamics (Dorri et al., 2018). Traditional multi-agent systems typically involve semi-autonomous agents coordinating through explicit protocols and structured messages to achieve common goals (Dunin-Keplicz & Verbrugge, 2011; Bakliwal et al., 2018). The advent of LLMs has revolutionized this landscape by enabling agents to utilize natural language for communication and collaboration in a believable proxy of human behavior (Park et al., 2023a), thereby fostering a more intuitive and flexible interaction model. Recent studies have further verified the superior performance of LLM-based multi-agent systems in various domains, such as programming, game playing, and complex reasoning tasks when compared to single-agent execution (Liu et al., 2023b; Wang et al., 2024; Light et al., 2023; Du et al., 2024). In this work, we strategically implement the power of LLM-based multi-agent systems to function as collaborative scientists, promoting de novo scientific ideas.
3 The Virtual Scientists
In this paper, we aim to build a multi-agent system using a real-world academic dataset to simulate how a scientist assembles a research team and collaboratively generates an abstract that details a novel scientific idea. Our VirSci system consists of two components: a scientific research ecosystem and a multi-agent system for scientific idea generation.
@@ 3.1 The Scientific Research Ecosystem
Past Paper Database To construct the past paper database $B_{\text{past}}$ using the Faiss, we selected papers published before the $y_{bound}$. Each paper includes essential information such as its title, citation count, and abstract.
@@
Contemporary Paper Database The contemporary paper database $B_{\text{con}}$, also constructed with Faiss, consists of papers published after $y_{bound}$. Similarly, each paper’s basic information is structured in the same way as the past papers. Although using papers from this time range may raise concerns about data leakage, given that LLMs are trained on data within this period, we will explain in detail why this does not pose a threat to the overall validity of our experiments and conclusions in the Appx. D.
@@
Author Knowledge Bank  For each scientist in $S$, we extract their basic profile from the computer science dataset, which includes their name, affiliations, citation count, research interests, and collaboration history. Using the KnowledgeBank module from AgentScope(Gao et al., 2024), we embed these scientist profiles into the author knowledge bank. This allows agents to quickly access and familiarize themselves with other initialized agents’ information. Notably, real author names are masked to prevent data leakage and privacy problems during agent initialization (See Appx. C).
@@
Adjacency Matrix Given the scientist set $S$, let $A$ represent the adjacency matrix, where $A_{i,j}$ denotes the number of times scientist $i$ has collaborated with scientist $j$. To prevent agents from always choosing previously collaborated scientists, overlooking the benefit of fresh collaborations that often lead to more original and impactful research (Zeng et al., 2021), we increment all values in $A$ by 1. This adjustment ensures that scientists with no prior collaborations still have a chance of being selected, encouraging agents to explore new partnerships.
3.2 The Multi-agent System for Scientific Collaboration
@@
We first randomly sample a scientist $s_{0}$ from $S$ as the team leader. The team leader then follows these steps to produce an abstract: (1) Collaborator Selection, (2) Topic Discussion, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. To help each agent become familiar with the backgrounds of other team members without overloading the initialization prompt, we employ retrieval-augmented generation (RAG) (Lewis et al., 2020), used throughout all five steps. All necessary prompts and example scenarios are shown in Appx. G and H.
@@
Collaborator Selection The first step in our system is collaborator selection, aimed at forming a team of scientists, $T=\{s_{0},\dots,s_{i},\dots,s_{n}\}$, where $n$ denotes the team size. When $s_{0}$ is selecting collaborators, we convert the adjacency matrix, $A$, into a probability distribution using the following equation: $P_{i,j}=\frac{A_{i,j}}{\sum_{j=1}^{N}A_{i,j}}$, where $N$ denotes the size of $S$. This allows the team leader to iteratively send invitations to preferred collaborators. Upon receiving an invitation, the invited scientist evaluates whether to join the team using the chain-of-thought process (Wei et al., 2022), considering the profiles of $s_{0}$ and the current team members. If accepted, the scientist is added to the team $T$. This process continues until the pre-defined team size $n$ is reached.
@@
Topic Discussion The next step is to propose a research topic, which will guide the research direction. Inspired by multi-round collaboration (Mezirow, 2003; Sunstein, 2005; Amgoud & Prade, 2009) and multi-agent collaboration strategies (Xu et al., 2023; Zhang et al., 2023; Shinn et al., 2024), we design a general team discussion mechanism. In this mechanism, team members engage in discussions based on a specific task description prompt. This process is applied not only to topic discussion but also to subsequent collaboration steps. While allowing agents to decide when to stop the discussion would better reflect real-world scenarios, fixing the number of turns ensures consistent inference costs across different team settings in our experiments. Therefore, we leave the discussion of adaptive turn numbers to the ablation study (See Sec. 5). Given the team $T$, the prompt during the topic discussion is:
@@
Idea Generation Third, the team is tasked with proposing several potential ideas. To align with genuine research workflows and mitigate LLM illusions (Huang et al., 2023), each agent is required to generate a comprehensive response that includes three key components: (1) a description of the idea, (2) a specific experimental plan, and (3) a self-assessment covering metrics such as novelty, feasibility, and clarity, representing the agent’s confidence (See Appnx. 12).
At the start of the idea generation process, when no ideas have yet been proposed, the agent is provided with references by searching $B_{past}$ using the topic $R_{topic}$, denoted as $B_{past}(R_{topic})$. The first idea-generation prompt is defined as:
$Q_{1,0}=\langle Q_{idea},R_{topic},B_{past}(R_{topic})\rangle,$ (2)
where $Q_{idea}$ represents the task description. Inspired by the concept of gradually expanding an archive of ideas (Zhang et al., 2023; Lu et al., 2024), when a scientist $s_{i}$ at turn $k$ receives an existing idea from the previous response $R_{k,i-1}$, we retain the previously generated ideas along with their corresponding references from $B_{past}$. These are passed to the next agent, who can either refine the existing idea or propose a new one, depending on its choice. The prompt is represented as:
$Q_{k,i}=\langle Q_{idea},R_{topic},B_{past}(R_{k,i-1}),\overset{k-1}{\underset{t=1}{\bigcup}}(\overline{D_{t}}),\overset{i-1}{\underset{j=0}{\bigcup}}(R_{k,j})\rangle.$ (3)
Afterwards, the response of $S_{i}$ at turn $k$ can be represented as $R_{k,i}\sim P_{s_{i}}(\cdot|Q_{k,i})$. After $K$ turns of discussion, we retain the three ideas with the highest confidence and store them in the idea list $I$.
Novelty Assessment To enhance the quality of ideas and mitigate agent overconfidence, we introduce an idea novelty assessment, enabling agents to compare each idea with related papers from $B_{past}$ and vote for the idea they consider most novel. Given the idea list $I$, agents search for related papers using each idea’s description to determine whether it significantly overlaps with existing works. To simulate a blind review process, no dialogue memory is included in the prompt. The prompt for $s_{i}$ at turn $k$ is defined as:
@@
where $I_{j}$ is the $j$-th idea in $I$. Following the chain-of-thought process, the response $R_{k,i}\sim P_{s_{i}}(\cdot|Q_{k,i})$ includes the scientist’s preferred idea and the reasoning behind their choice. The idea receiving the highest number of votes is then selected as the final idea, $R_{idea}$, for abstract generation.
Abstract Generation Lastly, the team is required to produce a comprehensive abstract that includes the following sections: (1) Introduction, (2) Objective, (3) Methods, (4) Expected Results, and (5) Conclusion (Alexandrov & Hennerici, 2007). At the start of abstract generation, the team leader provides an initial draft based on $R_{idea}$. The first abstract-generation prompt is:
$Q_{1,0}=\langle Q_{abstract},R_{idea}\rangle,$ (5)
where $Q_{abstract}$ represents the task description and format requirements.
When an abstract is provided by the previous response $R_{k,i-1}$, the next scientist’s response should include: (1) an evaluation of the prior abstract (evaluation metrics are detailed in Appx. 15), (2) proposed modifications, and (3) the revised abstract to enable continuous refinement. The corresponding prompt is:
$Q_{k,i}=\langle Q_{abstract},Q_{judgement},R_{k,i-1}\rangle,$ (6)
where $Q_{judgement}$ is the prompt that asks agents to evaluate the previous abstract. Dialogue history is not included in this prompt since the process is iterative and focuses on refining a single abstract. Including previous versions would make the prompt redundant. After $K$ turns of revision, the final abstract is denoted as $R_{abstract}$.
A self-review mechanism is also considered after $R_{abstract}$ is finalized to pre-check its novelty. The optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$, where more details are shown in Appx. E.1. Because it introduces uncertainty in total inference cost, making it difficult to ensure fair experimental comparisons, we only discuss the effectiveness of this module in the ablation study (see Sec. 4.4).

### v2

# Many Heads Are Better Than One: Improved Scientific Idea Generation by A LLM-Based Multi-Agent System

@@ Abstract
The rapid advancement of scientific progress requires innovative tools that can accelerate knowledge discovery. Although recent AI methods, particularly large language models (LLMs), have shown promise in tasks such as hypothesis generation and experimental design, they fall short of replicating the collaborative nature of real-world scientific practices, where diverse experts work together in teams to tackle complex problems. To address the limitations, we propose an LLM-based multi-agent system, i.e., Virtual Scientists (VirSci), designed to mimic the teamwork inherent in scientific research. VirSci organizes a team of agents to collaboratively generate, evaluate, and refine research ideas. Through comprehensive experiments, we demonstrate that this multi-agent approach outperforms the state-of-the-art method in producing novel scientific ideas. We further investigate the collaboration mechanisms that contribute to its tendency to produce ideas with higher novelty, offering valuable insights to guide future research and illuminating pathways toward building a robust system for autonomous scientific discovery. The code is available at https://github.com/open-sciencelab/Virtual-Scientists.
@@ 1 Introduction
The rapid acceleration of scientific research necessitates innovative tools for exploring new concepts and tackling complex challenges (park2023papers, ). Automatic scientific discovery has emerged as a promising solution to accelerate innovation, aligning with the ultimate goal within the scientific community (langley1987scientific, ). With the development of artificial intelligence (AI), automatic scientific discovery has the potential to revolutionize how research is conducted by automating key steps in the scientific process, ranging from hypothesis generation to experimental design (raghu2020survey, ; spangler2014automated, ).
Recent works like AI Scientist (lu2024ai, ), ResearchTown (yu2024researchtown, ), and HypoGen qi2024large leverage LLMs (openai, ; llama3, ) as intelligent agents to simulate the scientific idea generation process and advance automatic scientific discovery at various stages, including literature review (shi2023towards, ; hsu2024chime, ) and experimental design (wu2023autogen, ; huang2024crispr, ). However, these efforts either rely on a single-agent system, overlooking the collaborative nature of real-world research (kayacik2019identifying, ; gauch2003scientific, ; linsey2005collaborating, ), or employ an oversimplified collaboration framework and unrealistic data (e.g. personal information, collaboration networks) to model a multi-agent system, failing to capture real-world complexities. Consequently, they provide limited insights into multi-agent collaboration, which is essential for advancing autonomous scientific discovery.
@@
Figure 1: The proposed LLM-based multi-agent system, VirSci, includes five key steps: Collaborator Selection, where a research team is assembled; Topic Discussion, where the research topic is determined; Idea Generation, where team members propose and refine ideas; Novelty Assessment, where ideas are evaluated and voted on to select the best one; and Abstract Generation, where the selected idea is developed into a complete abstract.
To address these limitations, we build a virtual Scientific Research Ecosystem to benchmark the potential of multi-agent systems in scientific idea generation. Specifically, Scientific Research Ecosystem is a digit twin of research communities at the given time point, e.g., Jan 1st, 2024. It have three components: (1) 35,774,510 virtual scientists whose backgrounds and publications are cloned from real scientists at that moment, (2) a past paper database where papers are published before that moment and (3) a contemporary paper database where papers are published after that moment. To evaluate the novelty of the generated ideas, we introduce three metrics from different perspectives: dissimilarity to past papers, alignment with research trends, and the potential influence on contemporary research (shao2020bert, ; yang2022gender, ). By comparing the generated abstract against two paper databases, we ensure the generated ideas are innovative and align with emerging scientific directions, validating the effectiveness of our approach (Real examples and their professional analysis are shown in Appx. D).
Based on the virtual Scientific Research Ecosystems, we propose an LLM-based multi-agent system, Virtual Scientists (VirSci), designed to harness the potential of LLM agents in assisting autonomous scientific idea generation. Leveraging the inherent human-like reasoning capabilities of LLMs (xie2024can, ), VirSci simulates the collaborative process of generating scientific ideas (perry2017creativity, ; muzzio2024collaborative, ), by the following five steps (see Fig. 1): (1) Collaborator Selection, (2) Topic Selection, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. In Collaborator Selection stage, given a randomly selected agent as the team leader, it will exploit historical co-authors based on its collaboration history and academic social networks, while also exploring potential collaborators whose expertise and research interests align with the team’s goals march1991exploration ; rzhetsky2015choosing ; zeng2019increasing .
In Topic Selection stage, the scientists will discuss topics of common interest. The discussion will be terminated and restarted if a consensus cannot be achieved among the majority of the team. The scientists who are not interested can choose to leave the discussion at will. Otherwise, the discussion continues until a final topic is determined. In the Idea Generation stage, the virtual scientists retrieve relevant papers from the past paper database and engage in both inter- and intra-team discussions, where collaborators participate in iterative dialogues based on their backgrounds. This inter- and intra-team discussion distinguishes our approach from previous group discussion patterns zhang2023exploring ; qi2024large ; qian2024chat , enabling agents within the team to proactively seek advice from external agents (inter-team) through an “Invitation Mechanism”, while effectively balancing diverse perspectives within the team (intra-team). Once the discussions are over, the Abstract Generation stage begins, and the team generates a comprehensive abstract that encapsulates the proposed ideas.
We conduct extensive experiments to verify the effectiveness of VirSci in producing novel scientific ideas on both single-discipline and multi-discipline datasets. The findings prove that the multi-agent system improves the single-agent executive by, on average, +13.8% and +44.1% in alignment with and potential impacts on contemporary research, respectively. Furthermore, our experiments investigate collaboration mechanisms among agents that influence the performance of idea generation. The patterns observed in the experimental results align with findings from prior Science of Science studies (fortunato2018science, ; wu2019large, ; zeng2021fresh, ; shi2023surprising, ) published in Top venues, e.g., Science and Nature, providing valuable insights to guide future research toward autonomous scientific discovery.
Our core contributions are as follows:
@@
1) To the best of our knowledge, we propose the first multi-agent system with a scientific research ecosystem for conducting and benchmarking scientific collaborations, named VirSci, where real data is used for role-playing and objective evaluation.
2) To simulate a reliable scientific collaboration process, we propose an end-to-end pipeline that spans team organization to idea generation. A novel inter- and intra-team discussion mechanism is introduced to promote communication topology and enhance the simulation realism.
3) Extensive experiments demonstrate that multi-agent collaboration improves outcome quality, surpassing existing methods. We also conduct systematic experiments to investigate factors influencing the system, with findings aligning with established principles from the Science of Science, such as fresh teams tend to generate more innovative research, underscoring VirSci’s reliability as a powerful tool for autonomous scientific discovery.
@@ 2.1 AI for Scientific Discovery
In recent years, AI has transformed scientific discovery by offering powerful tools to enhance research processes (xu2021artificial, ). Generative AI, in particular, accelerates basic scientific discoveries by tackling complex tasks such as molecular identification (vignac2022digress, ), protein structure prediction (abramson2024accurate, ), and proteomics research (ding2024automating, ), significantly reducing experimental iteration times.
Besides, with the advent of LLMs, AI methodologies can step further and collaborate in streamlining critical stages of the scientific pipeline, including hypothesis generation, experimental design, data acquisition, and analysis(zheng2023large, ; wang2023scientific, ; miret2024llms, ; wysocki2024llm, ; lu2024ai, ; si2024can, ). Nevertheless, these approaches lack the collaborative nature of the scientists intrinsic to real-world research. VirSci is the first to harness the power of an LLM-based multi-agent system to facilitate the generation of research ideas in autonomous scientific discovery.
2.2 Collaboration in Multi-Agent Systems
A multi-agent system for team collaboration leverages autonomous agents to coordinate, communicate, and solve tasks within a shared environment, simulating human teamwork dynamics (dorri2018multi, ). Traditional systems typically rely on semi-autonomous agents using explicit protocols and structured messages to achieve shared goals (dunin2011teamwork, ; bakliwal2018multi, ). The emergence of LLMs has reshaped this landscape by enabling agents to communicate in natural language, fostering more intuitive and flexible interactions (park2023generative, ). Studies have demonstrated that LLM-based multi-agent systems outperform single-agent setups in tasks such as programming, game playing, and complex reasoning (du2024multi, ; wang2024rethinking, ; light2023avalonbench, ). However, prior multi-agent frameworks (qi2024large, ; yu2024researchtown, ) for focused scientific idea generation (1) lack dynamic team composition, restricting team members from accessing insights beyond their initial group, and (2) do not incorporate real scientist data and collaboration relationship, leading to impractical conclusions. To overcome these limitations, we introduce an inter- and intra-team discussion mechanism, enabling flexible collaboration among virtual scientists (supported by real data) and fostering the generation of de novo scientific ideas.
3 The VirSci
In this paper, we aim to build a multi-agent system using real-world academic datasets to simulate how a scientist assembles a research team and collaboratively generates an abstract that details a novel scientific idea [^3] An abstract effectively represents the key aspects of scientific research and serves as a concise reflection of its novelty. Additionally, considering the computational constraints, we focused our evaluation primarily on the generated abstracts rather than the full texts. This allows us to assess the system’s contributions while efficiently utilizing available resources. . Our VirSci system consists of two components: a scientific research ecosystem and a multi-agent system for scientific collaboration and idea generation.
@@ 3.1 The Scientific Research Ecosystem
Past Paper Database. To construct the past paper database $B_{\text{past}}$ using the Faiss [^4] Faiss is a Python library designed for efficient similarity search and clustering of dense vectors. johnson2019billion , we selected papers published before the $y_{bound}$. Each paper includes essential information such as its title, citation count, and abstract.
@@
Contemporary Paper Database. The contemporary paper database $B_{\text{con}}$, also constructed with Faiss, consists of papers published after $y_{bound}$. Similarly, each paper’s basic information is structured in the same way as the past papers. Although using papers from this time range may raise concerns about data leakage, given that LLMs are trained on data within this period, we will explain in detail why this does not pose a threat to the overall validity of our experiments in Appx. A.
@@
Author Knowledge Bank. For each scientist in $S$, we extract their basic profile from the pre-processed dataset (More details are shown in Sec. 4.1), which includes their name, affiliations, citation count, research interests, and collaboration history. Using the KnowledgeBank module from AgentScope (agentscope, ), we embed these scientist profiles into the author knowledge bank. This allows agents to quickly access and familiarize themselves with other initialized agents’ information. Notably, real author names are masked to prevent data leakage and privacy problems during agent initialization.
@@
Adjacency Matrix. Given the scientist set $S$, let $A$ represent the adjacency matrix, where $A_{i,j}$ denotes the number of times that scientist $i$ has collaborated with scientist $j$. To prevent agents from always choosing previously collaborated scientists, overlooking the benefit of fresh collaborations that often lead to more original and impactful research (zeng2021fresh, ), we increment all values in $A$ by 1 (more experiments about the increment function in Sec. 5.4). This adjustment ensures an explore-exploit mechanism, allowing scientists with no prior collaborations a chance to be selected, thereby encouraging agents to explore new partnerships.
3.2 The Multi-agent System
@@
We first randomly sample a scientist $s_{0}$ from $S$ as the team leader. The team leader then follows these steps to produce an abstract: (1) Collaborator Selection, (2) Topic Discussion, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. To help each agent become familiar with the backgrounds of other team members without overloading the initialization prompt, we employ retrieval-augmented generation (RAG) (lewis2020retrieval, ), used throughout all five steps. All necessary prompts and example scenarios are shown in Appx. E and F.
@@
Collaborator Selection. The first step in our system is to form a team of scientists, $T=\{s_{0},\dots,s_{i},\dots,s_{n}\}$, where $n$ denotes the team size. When $s_{0}$ is selecting collaborators, we convert the adjacency matrix, $A$, into a probability distribution using the following equation: $P_{i,j}=\frac{A_{i,j}}{\sum_{j=1}^{N}A_{i,j}}$, where $N$ denotes the size of $S$. This allows the team leader to iteratively send invitations to preferred collaborators. Upon receiving an invitation, the invited scientist evaluates whether to join the team using the chain-of-thought process (wei2022chain, ), considering the profiles of $s_{0}$ and the current team members. If accepted, the scientist is added to the team $T$. This process continues until the pre-defined team size $n$ is reached.
@@
Topic Discussion. The next step is to propose a research topic, which will guide the research direction. Inspired by multi-round collaboration (mezirow2003critical, ; sunstein2005societies, ; amgoud2009using, ) and multi-agent collaboration strategies (xu2023magic, ; zhang2023exploring, ; shinn2024reflexion, ), we design an inter- and intra-team discussion mechanism. In this mechanism, team members engage in discussions based on a specific task description prompt (inter-team discussion). The default discussion mechanism between agents follows a round-table format with a sequential progression (Additional discussion topologies are presented in Sec. 5.3.1). This process is also applied to subsequent collaboration steps. While allowing agents to decide when to stop the discussion would better reflect real-world scenarios, fixing the number of turns ensures consistent inference costs across different team settings in our experiments. Therefore, we leave the discussion of adaptive turn numbers to the ablation study (See Sec. 5.3.2). The prompt for agent $i$ during the topic discussion is:
@@
Idea Generation. Third, the team is tasked with proposing several potential ideas. To align with genuine research workflows and mitigate LLM illusions (huang2023survey, ), each agent is required to generate a comprehensive response that includes three key components: (1) a description of the idea, (2) a specific experimental plan, and (3) a self-assessment covering metrics such as novelty, feasibility, and clarity, representing the agent’s confidence (See Appx. Fig. 20).
At the start of the idea generation process, when no ideas have yet been proposed, the agent is provided with references by searching $B_{past}$ using the topic $R_{topic}$, denoted as $B_{past}(R_{topic})$. The first prompt is defined as: $Q_{1,0}=\langle Q_{idea},R_{topic},B_{past}(R_{topic})\rangle$, where $Q_{idea}$ represents the task description.
Inspired by the concept of gradually expanding an archive of ideas (zhang2023exploring, ; lu2024ai, ), when a scientist $s_{i}$ at turn $k$ receives an existing idea from the previous response $R_{k,i-1}$, we retain the previously generated ideas along with their corresponding references from $B_{past}$. These are passed to the next agent, who can either refine the existing idea or propose a new one, depending on its choice. The prompt $Q_{k,i}$ is represented as:
$\langle Q_{idea},R_{topic},B_{past}(R_{k,i-1}),\overset{k-1}{\underset{t=1}{\bigcup}}(\overline{D_{t}}),\overset{i-1}{\underset{j=0}{\bigcup}}(R_{k,j})\rangle.$ (2)
Afterwards, the response of $S_{i}$ at turn $k$ can be represented as $R_{k,i}\sim P_{s_{i}}(\cdot\mid Q_{k,i})$. After $K$ turns of discussion, we retain the three ideas with the highest confidence and store them in the idea list $I$.
Novelty Assessment. To enhance the quality of ideas and mitigate agent overconfidence, we introduce an idea novelty assessment, enabling agents to compare each idea with related papers from $B_{past}$ and vote for the idea they consider most novel. Given the idea list $I$, agents search for related papers using each idea’s description to determine whether it significantly overlaps with existing works. To simulate a blind review process, no dialogue memory is included in the prompt. The prompt for $s_{i}$ at turn $k$ is defined as:
@@
where $I_{j}$ is the $j$-th idea in $I$. Following chain-of-thought process, the response $R_{k,i}\sim P_{s_{i}}(\cdot\mid Q_{k,i})$ includes the scientist’s preferred idea and the reasoning behind their choice. The idea receiving the highest votes is then selected as the final idea, $R_{idea}$, for abstract generation. The effectiveness of novelty assessment is discussed in Sec. 5.2.
Abstract Generation. Lastly, the team is required to produce a comprehensive abstract that includes the following sections: (1) Introduction, (2) Objective, (3) Methods, (4) Expected Results, and (5) Conclusion (alexandrov2007writing, ). At the start of abstract generation, the team leader provides an initial draft based on $R_{idea}$. The first abstract-generation prompt is: $Q_{1,0}=\langle Q_{abstract},R_{idea}\rangle$, where $Q_{abstract}$ represents the task description and format requirements.
When an abstract is provided by the previous response $R_{k,i-1}$, the next scientist’s response should include: (1) an evaluation of the prior abstract (evaluation metrics are detailed in Appx. Fig. 23), (2) proposed modifications, and (3) the revised abstract to enable continuous refinement. The prompt is defined as:
$Q_{k,i}=\langle Q_{abstract},Q_{judgment},R_{k,i-1}\rangle,$ (4)
where $Q_{judgment}$ is the prompt that asks agents to evaluate the previous abstract. Dialogue history is not included in this prompt since the process is iterative and focuses on refining a single abstract. Including previous versions would make the prompt redundant. After $K$ turns of revision, the final abstract is denoted as $R_{abstract}$.
A self-review mechanism is also considered after $R_{abstract}$ is finalized to pre-check its novelty. In this self-review, the optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$. The prompt is:
$\begin{split}Q_{1,0}=\langle Q_{abstract},Q_{judgement},R_{review},\\ B_{past}(R_{abstract}),R_{abstract}\rangle\end{split}$ (6)